OpenStudy (anonymous):

The velocity of a car decelerates from 30ms-1 to 15ms-1 within a distance of 75m. Calculate the extra distance travelled by the car before it stops.

OpenStudy (anonymous):

a=(v^2-u^2)/2s with v=15 u=30 find a and then use S=(v^2-u^2)/2a with v=0 u=15